Solve the system by the elimination method.
5x + 5y = -13
7x - 3y = 19
I'm having a hard time understading this equation.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Solve the system by the elimination method.
5x + 5y = -13
7x - 3y = 19
--------------------------
Multiply thru the 1st Eq by 3
Multiply thru the 2nd Eq by 5
--------------------------
15x + 15y = -39
35x - 15y = 95
------------
Add and solve for "x":
50x = 46
x = 23/25
-------------------
Substitute into 5x + 4y = -13
5(23/25) + 4y = -13
(23/5) + 4y = (-65/5)
4y = -88/5
y = -22/5
=====================
Cheers,
Stan H.

Answer by Alan3354(69443)

Solve the system by the elimination method.
5x + 5y = -13
7x - 3y = 19
-------------
To eliminate one of the variables, multiply one or both eqns to get the same coefficient for one of them. You can do either, I'll do y
--------
5x + 5y = -13 x3 --> 15x + 15y = -39
7x - 3y = 19 x5 --> 35x - 15y = 95
--------------------------------------
15x + 15y = -39
35x - 15y = 95
--------------- Add
50x = 56
x = 56/50 = 1.12
-----------------
Usually it's "elimination and substution", once you find one, you can sub and find the others.
7x - 3y = 19
7*1.12 - 3y = 19
3y = 7.84 - 19 = - 11.16
y = -3.72